# SMC++: Masked Learning of Unsupervised Video Semantic Compression

### v1

@@ 1 Introduction
To tackle this problem, lots of research efforts have been devoted to emphasizing the coding of semantics-relevant information. For example, early works [7][8][9] and standards [10][11][12] additionally transport the manually-designed image descriptors. Later, some methods [13][14][15] improve the traditional codec with hand-crafted designs to better cope with the specific tasks. Meanwhile, some methods [16][17][18][19][20][21] compress the feature maps of AI models instead of the images, where the downstream task modules shall be fine-tuned to accommodate the features. Besides, scalable coding methods [22][23][24][25][26][27][28] divide the bitstream into two parts, i.e., one for analysis tasks and another one for video reconstruction. The former part is usually optimized by task-specific losses.
@@
This work extends our preliminary conference version [36] with the following substantial improvements. 1) We provide more analysis and visualization results of the masked learning procedure and the proposed NSS strategy, aiming to help readers more intuitively understand their effectiveness in the context of video semantic compression. 2) We introduce an additional masked motion prediction term to the original appearance-only MVM task loss. We compare and analyze the impact of different motion target choices. 3) We propose a Blueprint-guided compression Transformer (Blue-Tr) for more effective semantic compression. We perform ablation studies and detailed analysis for our proposed Blue-Tr. 4) The above two improvements lead to the advanced SMC++ model, which remarkably outperforms the most recent semantic coding approach VCS [37] and neural video codec DCVC-FM [38]. We also compare recent approaches under a more advanced compression setting, i.e., group-of-picture(GOP) size of 32. 5) We append the SMC++ model with a lightweight decoder, obtaining a PSNR-oriented model SMC++*. SMC++* surpasses the advanced VTM20.0 software [39] by a notable margin, in both low-delay P (LDP) and random access (RA) test conditions.
The rest of the paper is organized as follows. In Section 2, we give a brief review of the related works. In Section 3, we outline our masked learning-based video semantic coding framework, then instantiate it as two specific models. Finally, we report the experimental results and conclude the paper in Section 4 and Section 5, respectively.
@@ 2 Related Works
Video Compression. Previous video codecs, including traditional ones [40][1][2][41], learnable ones [42][3][4][43][44][45][46][42][47][48][38] and mixed ones [49][50][51], are designed to achieve better pixel-wise signal quality metrics, e.g., PSNR and MS-SSIM [52], which mainly serve the human visual experience. Recently, there are also some generative video coding methods [53][54] that mainly consider visual comfort and perceptual quality. Despite recent methods achieving significant strides in low-level visual metrics like PSNR and SSIM, their effectiveness on semantic AI tasks is still undesirable [6]. This prompts the research on the video semantic compression problem.
Video Coding for Machine (VCM). Early standards such as CDVA [12] and CDVS [11][10] propose to pre-extract and transport the image keypoints, supporting image indexing or retrieval tasks. Some works[7][8][9][16][17][18][19][20] compress the intermediate feature maps instead of images. Besides, some works [13][14][15][55] [56][57][58] improve traditional codecs by introducing downstream task-guided rate-distortion optimization strategy or another task-specific feature encoding stream [59][8]. Also, some methods [60][61][62][63][26] optimize the learnable codecs by directly incorporating the downstream task loss. Recently, some methods exploit hand-crafted structure maps [64][65] for semantic coding. To summarize, most above methods rely on task-specific labels or hand-crafted priors, which may not be generalizable enough.
Recently, some works have leveraged self-supervised representation learning methods for learning a compact semantic representation. As a pioneering work, Dubois et al.[66] theoretically revealed that the distortion term of the lossy rate-distortion trade-off for image classification can be approximated by a contrastive learning objective [67]. However, the compressed semantics are empirically effective to a group of tasks that share a similar prior (i.e. , labels are invariant to data augmentations), but may severely discard the semantics useful to other tasks such as detection and segmentation. Feng et al. [20] proposed to learn a unified feature representation for AI tasks from unlabeled data in a similar manner. In these methods, the downstream models are required to be fine-tuned to adapt to the features.
@@
Scalable Coding and Visual-Semantic Fusion Coding. Scalable coding methods [64][23][22][24][25][56] can achieve excellent compression efficiency when measured with the trained tasks, but usually show undesirable results on the tasks/data out of the training scope, due to the supervised learning paradigm. Visual-semantic fusion coding methods [65][68][69][64][70], a.k.a, conceptual coding [71][72][73][74], first extract the structure information and the texture information on the encoding side, and then fuse the two parts into a image on the decoding side. The fused images are readily fed into various tasks and achieve superior performance even at low bitrate levels. However, almost all these methods employ semantic segmentation map [65] or edge map [64] as the semantic stream, not fully discarding the task-specific priors.
Compressed Video Analysis. There are also amounts of works such as [75][76][77][78][79][80] perform video analysis tasks, such as action recognition [81][82][83][84][85][86] and multiple object tracking (MOT) [87][88], in the compressed video domain. However, these methods focus on developing video analysis models that better leverage the partially decoded video stream, such as the motion vector. In contrast, our work focuses on the coding procedure.
Self-Supervised Semantic Learning. Recent methods can be mainly divided into two catogories, i.e., Contrative Learning (CL) ones  [67][89][90][91] and Masked Image Modeling (MIM) ones [92][31][32][33]. CL methods use two augmented views of the same image as the positive pair, and other images as negative samples. The learned semantics relies on the employed augmentation strategy [93], and is usually biased to global semantics. Recently, MIM methods have gained increasing attention. MIM simply predicts masked patches from unmasked ones, while showing remarkably strong performance in downstream tasks. After the pioneering works, e.g., MAE [31] and Beit [94], amounts of works are proposed to improve the MIM framework [95][96][33] or the prediction target [97][92][98].
Although MIM methods are superior to CL ones in many aspects, when the masked region reconstruction loss serves as the semantic learning objective of a compression system, it also facilitates the system encoding some non-semantic information, and wasting extra bitcosts. Our work solves this problem by explicitly suppressing the non-semantics information within the MAE feature space. Very recently, there are some works [99][100] extending VideoMAE [32] with motion prediction targets, achieving superior results on video action recognition task. In this work, we also incorporate this motion-based target into our system, verifying its effectiveness on the video semantic compression problem.
@@ 3.1 Framework Overview
Semantic Extraction Network (Sem-Net). To transform the videos from RGB space to semantic space, the original video $X$ and the lossy video $\tilde{X}$ are encoded as the semantic features $S$ and $\tilde{S}$, respectively. The tensor shape of $S$ and $\hat{S}$ are both $\mathbb{R}^{T\times 512\times\frac{H}{32}\times\frac{W}{32}}$, where $T$ and $H\times W$ represent the temporal length and spatial dimensions of the input video. For producing $S$, we adopt the ResNet18 [101] network as the Sem-Net, but replace its first Max-pooling layer with a stride two convolution layer for retaining more information. For producing $\tilde{S}$, we adopt a more lightweight network denoted Sem-Nets, since it has to be performed twice on both the encoder and decoder sides. Sem-Nets mainly consists of five convolution layers of stride size two and kernel size three. We also insert depth-wise convolutions of large kernel size 7$\times$7 to effectively enlarge the receptive field size, without substantially increasing the computational cost. The weights of the above two networks are randomly initialized.
@@
Semantic-Visual Information Fusion. After obtaining the reconstructed semantic feature $\hat{S}$ and the lossy video $\tilde{X}$ on the decoder side, we use a UNet-style [102] generator network termed F-Net to synthesize the final video $\hat{X}$, which will be consumed by downstream AI task models. Considering that global structures instead of local details are more critical to video semantics, the above feature fusion procedures within the UNet are conducted in four-, eight-, and sixteen-times downsampled feature spaces.
@@ 3.2 Semantic Learning Objective
MAE Learning. Given the decoded video $\hat{X}$, we first divide it into regular non-overlapping patches of size $16\times 16$, and each patch is transformed to tokens by linear embedding. Then, a proportion of tokens are randomly masked, and the remaining unmasked token set ${Tok}$ is fed into the prediction network $\phi$, which includes an encoder and a decoder, for reconstructing the video. The pixel-wise reconstruction loss of the MAE task is given by,
@@
where $i$ is the masked token index, $\mathcal{M}$ is the index set of masked tokens, and $X$ is the ground-truth video, ||$\cdot$|| ⋅ | denotes the $\ell$1 distance function. We adopt the $\ell$1 instead of the patch-normalized $\ell$[31][32], as we prefer the stable gradient values 1 of $\ell$1. This stabilizes the joint training with other loss functions in our system, such as GAN loss.
@@
$\mathcal{L}_{MAE-motion}=\frac{1}{\mathcal{M}}\sum_{i\in\mathcal{M}}|\mathcal{T}\cdot\phi_{m}(Tok)[i]-\mathcal{T}\cdot{X}[i]|,$ (2)
where $\phi_{m}$ shares the parameters of the encoder with $\phi$ but includes a motion decoder that is not shared. $\mathcal{T}$ denotes the motion calculation function, such as the pixel-wise difference map, the optical flow map, and the feature-wise correlation map. In the experimental section, we extensively compare these design choices.
@@
Non-Semantics Suppressed (NSS) Learning. To suppress the non-semantic information leaked from $X$ to the compressed video $\hat{X}$, we explicitly regularize the information entropy of the tokens $Tok$ in MAE feature space, conditioned on the mined semantic feature $S$. However, due to the difficulty of estimating the entropy of a continuous variable, we insert a quantization operation in the tail of the tokenization procedure, so that $Tok$ is a discrete variable. Then, we use a Gaussian Mixture Model (GMM) [103] with component number $K$ to approximate its distribution. The distribution of each token ${Tok}(i)$ is defined by the dynamic mixture weights $w_{i}$, means $\mu_{i}$ and log variances $\sigma_{i}$, which are produced by a density parameter estimation network (Density-Net). With these parameters, the distributions can be determined as,
@@
$p(Tok[i]|{S})=c({Tok}[i]+0.5)-c(Tok[i]-0.5),$ (4)
where $c(\cdot)$ is the cumulative function [104] of the GMM in Equation 3. Finally, the non-semantics suppressed mask learning objective can be given by,
$\mathcal{L}_{Sem}=-\beta\log(p({Tok}|{S}))+\mathcal{L}_{MAE},$ (5)
where $\beta$ is the balancing weight.
@@
Discussion. Although Equation 4 shares a similar format with the bit estimation procedure of the hyper-prior-based image compression methods [105][106], our goal is fundamentally different from theirs. Our approach aims to suppress the extra non-semantic information that is introduced by the MAE task, transporting zero bits, while the method [105] explores the hierarchical redundancies within images, and the estimated bits are additionally transported.
Refer to caption Refer to caption
(a) (b)
We further visualize the patch features, aiming to intuitively illustrate how the entropy regularization item of NSS improves the effectiveness of the semantics learned by MAE. First, we randomly select one hundred videos from the UCF101 dataset. Then, we cluster the patch features of these videos into individual groups using the K-Means algorithm [107]. The cluster ID is then utilized as the pseudo-category for each patch. Subsequently, the features are visualized using the t-SNE technique [108]. As compared in Figure 2, although the vanilla MAE loss can learn roughly separable semantic representation, our non-semantics suppressed loss $\mathcal{L}_{Sem}$ learns much clearer semantic boundaries, indicating less semantic noises. This also saves the semantic bitcost (0.0074bpp v.s. 0.0028bpp for $\mathcal{L}_{MAE}$ and $\mathcal{L}_{Sem}$, respectively).
Density-Net. It consists of two stacked convolutions of kernel size three, aligning the semantic feature ${S}$ to the MAE feature space. Then, we append a multiple layer perceptron (MLP) to predict $w$, $\mu$, and $\sigma$ within Equation 3. The gradient of ${S}$ is detached during the back-propagation procedure, forming a self-bootstrapping paradigm and avoiding the collapsing solution [109]. NSS scheme enforces the masked learning procedure to focus on mining semantic information, while the optimized semantic-rich ${S}$ leads to a more principled NSS objective.

### v2

@@ 1 Introduction
To tackle this problem, lots of research efforts have been devoted to emphasizing the coding of semantics-relevant information. For example, early works [7][8][9] and standards [10][11][12] additionally transport the manually-designed image descriptors. Later, some methods[13][14] improve the traditional codec with hand-crafted designs to better cope with the specific tasks. Meanwhile, some methods [15][16][17][18][19][20] compress the feature maps of AI models instead of the images, where the downstream task modules shall be fine-tuned to accommodate the features. Besides, scalable coding methods [21][22][23][24][25][26][27] divide the bitstream into two parts, i.e., one for analysis tasks and another one for video reconstruction. The former part is usually optimized by task-specific losses.
@@
This work extends our preliminary conference version [35] with the following substantial improvements. 1) We theoretically demonstrate that optimizing the NSS loss is equivalent to maximizing the mutual information between the compressed video patch tokens and their corresponding semantic labels, thereby improving semantic separability. In addition, we introduce an extra masked motion prediction term to the original appearance-only MVM task loss. We also offer further analysis and visualization of the masked learning procedure in the supplementary material. 2) We propose a Blueprint-guided compression Transformer (Blue-Tr) to achieve more effective semantic compression, upgrading the SMC model to the SMC++ model. By appending it with a lightweight detail decoder, we obtain the PSNR-oriented SMC++* model. 3) We conduct more comprehensive evaluations, including results on video question answering and fine-grained action recognition tasks, results on LLM-based downstream models, the flexibility of using multiple codecs (VVC, H.264, H.265, and DCVC-FM) as the base codec, and comparisons with more recent methods. Ablation studies are performed on MOT and VOS tasks, on hyperparameter selection, teacher-model-based MAE learning, various loss terms, attention mechanisms, the number of historical frames used as context, etc.
@@ 2 Related Works
Video Compression. Previous video codecs, including traditional ones [36][1][2] and learnable ones [3][4][37][38][39][40][41][42][43], are designed to achieve better pixel-wise signal quality metrics, e.g., PSNR and MS-SSIM [44], which mainly serve the human visual experience. Recently, there are also some generative video coding methods [45][46] that mainly consider visual comfort and perceptual quality. Despite recent methods achieving significant strides in low-level visual metrics like PSNR and SSIM, their effectiveness on semantic AI tasks is still undesirable [6][47][48][49][50][51][52]. This prompts the research on the video semantic compression problem.
Video Coding for Machine (VCM). Early standards such as CDVA [12] and CDVS [11] propose to pre-extract and transport the image keypoints, supporting image indexing or retrieval tasks. Some works [15][16][17][18][19] compress the intermediate feature maps instead of images. Besides, some works [14][53] [54][55] improve traditional codecs by introducing downstream task-guided rate-distortion optimization strategy or another task-specific feature encoding stream [56][8]. Also, some methods [57][25] optimize the learnable codecs by directly incorporating the downstream task loss. Recently, some methods exploit hand-crafted structure maps [58][59] for semantic coding. Most above methods rely on task-specific labels or hand-crafted priors, which may not be generalizable enough.
Recently, some works have leveraged self-supervised representation learning methods for learning a compact semantic representation. As a pioneering work, Dubois et al.[60] theoretically revealed that the distortion term of the lossy rate-distortion trade-off for image classification can be approximated by a contrastive learning objective [61]. However, the compressed semantics are empirically effective to a group of tasks that share a similar prior (i.e. , labels are invariant to data augmentations), but may severely discard the semantics useful to other tasks such as detection, segmentation, and even medical tasks [62][63][64][65]. Feng et al. [19][66][67] proposed to learn a unified feature representation for AI tasks from unlabeled data in a similar manner. In these methods, the downstream models are required to be fine-tuned to adapt to the features.
As for the incorporation of motion information, most previous approaches [68][69] directly adopt the architecture of PSNR-oriented codecs [37][4], utilizing motion information such as optical flow to compress inter-frame redundancy. CMVC [70] integrates motion to guide a diffusion model for generating smooth and natural videos, in ultra-low-bitrate video generative compression. HMFVC [71] extracts semantic information from motion flow for machine analysis. However, none of the previous approaches utilize motion as a learning target, to facilitate preserving temporal semantics during compression.
@@
Scalable Coding and Visual-Semantic Fusion Coding. Scalable coding methods [58][22][21][23][24][73] can achieve excellent compression efficiency when measured with the trained tasks, but usually show undesirable results on the tasks/data out of the training scope, due to the supervised learning paradigm. Visual-semantic fusion coding methods [59][74][75][58], a.k.a, conceptual coding [76][77][78][79], first extract the structure information and the texture information on the encoding side, and then fuse the two parts into a image on the decoding side. The fused images are readily fed into various tasks and achieve superior performance even at low bitrate levels. However, almost all these methods employ semantic segmentation map [59] or edge map [58] as the semantic stream, not fully discarding the task-specific priors.
Compressed Video Analysis. There are also amounts of works such as [80][81] perform video analysis tasks, such as action recognition [82][83][84][85][86] and multiple object tracking (MOT) [87], in the compressed video domain. However, these methods focus on developing video analysis models that better leverage the partially decoded video stream, such as the motion vector. In contrast, our work focuses on the coding procedure.
Self-Supervised Semantic Learning. Recent methods can be mainly divided into two catogories, i.e., Contrative Learning (CL) ones  [61][88] and Masked Auto-Encoder (MAE) learning ones [30][31]. MAE simply predicts masked patches from unmasked ones, while showing remarkably strong performance in downstream tasks. Our work adapts the MAE to better address the compression problem, by explicitly suppressing non-semantic information within its feature space.
Motion-Enhanced MAE. There are also some works that utilize motion information as the MAE target, aiming to capture more temporal semantics in videos. MGMAE [89] leverages optical flow to build a temporally-consistent masking strategy, reducing information leakage across frames. Masked Motion Encoding [90] and MotionMAE [91] and Song et al. [92] extend the reconstruction task by not only predicting static appearance, but also motion trajectories from masked regions. MASA [93] introduces a motion-aware masked autoencoder for sign language recognition. Our work is the first to introduce the masked motion reconstruction target to the field of video compression.
@@ 3.1 Framework Overview
Semantic Extraction Network (Sem-Net). To transform the videos from RGB space to semantic space, the original video $X$ and the lossy video $\tilde{X}$ are encoded as the semantic features $S$ and $\tilde{S}$, respectively. The tensor shape of $S$ and $\hat{S}$ are both $\mathbb{R}^{T\times 512\times\frac{H}{32}\times\frac{W}{32}}$, where $T$ and $H\times W$ represent the temporal length and spatial dimensions of the input video. For producing $S$, we adopt the ResNet18 [94] network as the Sem-Net, but replace its first Max-pooling layer with a stride two convolution layer for retaining more information. For producing $\tilde{S}$, we adopt a more lightweight network denoted Sem-Nets, since it has to be performed twice on both the encoder and decoder sides. Sem-Nets mainly consists of five convolution layers of stride size two and kernel size three. We also insert depth-wise convolutions of large kernel size $\times7$ to effectively enlarge the receptive field size, without substantially increasing the computational cost. The weights of the above two networks are randomly initialized.
@@
Semantic-Visual Information Fusion. After obtaining the reconstructed semantic feature $\hat{S}$ and the lossy video $\tilde{X}$ on the decoder side, we use a UNet-style generator network termed F-Net to synthesize the final video $\hat{X}$, which will be consumed by downstream AI task models. Considering that global structures instead of local details are more critical to video semantics, the above feature fusion procedures within the UNet are conducted in four-, eight-, and sixteen-times downsampled feature spaces.
@@
Discussion with the Scalable Coding Approaches [21][23]. These approaches [21, 23] use the base layer to deliver semantic information, whereas our approach uses the enhancement layer for this purpose. This provides the following benefits: First, the base-layer semantic information in [21, 23] is task-specific. This contradicts the core principle of our framework: “A single model that generates generalizable bitstream for various tasks”. In contrast, we utilize a traditional codec as the base layer, which is task-agnostic and not bound to specific tasks. Second, the bitstream produced by our method is partially compatible with traditional codec decoders. This compatibility allows the video stream to meet diverse user requirements. For example, users with low-end devices incapable of running deep learning models can still watch video, using the traditional codec for decoding.
@@ 3.2 Semantic Learning Objective
MAE Learning. Given the decoded video $\hat{X}$, we first divide it into regular non-overlapping patches of size $16\times 16$, and each patch is transformed to tokens by linear embedding. Then, a proportion of tokens are randomly masked, and the remaining unmasked token set Tok{Tok} is fed into the prediction network $\phi$, which includes an encoder and a decoder, for reconstructing the video. The pixel-wise reconstruction loss of the MAE task is given by,
@@
where $i$ is the masked token index, $\mathcal{M}$ is the index set of masked tokens, and $X$ is the ground-truth video, $|\cdot|$ denotes the $\ell_{1}$ distance function. We adopt the $\ell_{1}$ instead of the patch-normalized $\ell_{2}$ [30][31], as we prefer the stable gradient values 1 of $\ell_{1}$. This stabilizes the joint training with other loss functions in our system, such as GAN loss.
@@
$\mathcal{L}_{MAE-motion}=\frac{1}{\mathcal{M}}\sum_{i\in\mathcal{M}}|\phi_{m}(Tok)[i]-\mathcal{T}\cdot{X}[i]|,$ (2)
where $\phi_{m}$ shares the parameters of the encoder with $\phi$ but includes a motion decoder that is not shared. Compared to the appearance loss in the vanilla MAE framework, the motion loss ensures that motion pattern information is also preserved in the token space. This enhancement benefits a wide range of video-related tasks by capturing both spatial and temporal dynamics effectively. $\mathcal{T}$ denotes the motion calculation function, such as the RGB difference map and the optical flow map. In the experimental section, we extensively compare these design choices.
@@
Non-Semantics Suppressed (NSS) Semantic Learning. To suppress the non-semantic information leaked from $X$ to the compressed video $\hat{X}$, we explicitly regularize the information entropy of the tokens $Tok$ in MAE feature space, conditioned on the mined semantic feature $S$. However, due to the difficulty of estimating the entropy of a continuous variable, we insert a quantization operation in the tail of the tokenization procedure, so that $Tok$ is a discrete variable. Then, we use a Gaussian Mixture Model (GMM) [95] with component number $K$ to approximate its distribution. The distribution of each token Tok(i){Tok}(i) is defined by the dynamic mixture weights $w_{i}$, means $\mu_{i}$ and log variances $\sigma_{i}$, which are produced by a density parameter estimation network (Density-Net). With these parameters, the distributions can be determined as,
@@
p(Tok[i]|S)=c(Tok[i]+0.5)c(Tok[i]0.5),p(Tok[i]|{S})=c({Tok}[i]+0.5)-c(Tok[i]-0.5), (4)
where $c(\cdot)$ is the cumulative function [96] of the GMM in Equation (3). Finally, the non-semantics suppressed mask learning objective can be given by,
$\mathcal{L}_{Sem}=-\beta\frac{1}{\mathcal{M}}\sum_{i\in\mathcal{M}}\log(p({Tok[i]}\mid{S}))+\mathcal{L}_{MAE},$ (5)
where $\mathcal{M}$ denotes the index set of masked tokens. $\beta$ is the balancing weight.
@@
Discussion. Although Equation (4) shares a similar format with the bit estimation procedure of the hyper-prior-based image compression methods [97][98], our goal is fundamentally different from theirs. Our approach aims to suppress the extra non-semantic information that is introduced by the MAE task, transporting zero bits, while the method [97] explores the hierarchical redundancies within images, and the estimated bits are additionally transported.
Density-Net. It consists of two stacked convolutions of kernel size three, aligning the semantic feature S{S} to the MAE feature space. Then, we append a multiple layer perceptron (MLP) to predict $w$, $\mu$, and $\sigma$ within Equation (3). The gradient of S{S} is detached during the back-propagation procedure, forming a self-bootstrapping paradigm and avoiding the collapsing solution [99]. NSS scheme enforces the masked learning procedure to focus on mining semantic information, while the optimized semantic-rich S{S} leads to a more principled NSS objective.
@@ 3.3 Overall Optimization Target
3.4 Theoretical Justification of the NSS Objective
Notation. To begin, we introduce the notations used throughout this section. We refer to video patch token features and the semantic features, of all videos, as the random variables $\mathbf{{TOK}}$ and $\mathbf{{S}}$. The true conditional distribution of $\mathbf{TOK}$ given $\mathbf{S}$ is denoted as $P_{r}(\mathbf{TOK}\mid\mathbf{S})$, which will be simplified to $P_{r}$ in the following for clarity. The distribution estimated by Density-Net is represented as $P(\mathbf{TOK}\mid\mathbf{S})$. Let $\mathcal{C}$ be the discrete representation of $\mathbf{S}$ obtained via a clustering algorithm such as KMeans [103].
Lemma 1: Upper Bound on Conditional Entropy. To establish the theoretical basis, we begin by analyzing the KL divergence between $P_{r}$ and $P(\mathbf{TOK}\mid\mathbf{S})$,
$KL\Bigl(P_{r}\,\|\,P\Bigr)=\mathbb{E}_{P_{r}}\!\left[\log\frac{P_{r}(\mathbf{TOK}\mid\mathbf{S})}{P(\mathbf{TOK}\mid\mathbf{S})}\right]\geq 0.$ (7)
Expanding the logarithm yields,
$\mathbb{E}_{P_{r}}\!\left[\log P_{r}(\mathbf{TOK}\mid\mathbf{S})-\log P(\mathbf{TOK}\mid\mathbf{S})\right]\geq 0,$ (8)
then,
$\mathbb{E}_{P_{r}}\,\log P(\mathbf{TOK}\mid\mathbf{S})\geq-\mathbb{E}_{P_{r}}\,\log P_{r}\big(\mathbf{TOK}|\mathbf{S}\big),$ (9)
where
$\mathbb{E}_{P_{r}}\!\left[\log P_{r}\big(\mathbf{TOK}|\mathbf{S}\big)\right]=H\big(\mathbf{TOK}|\mathbf{S}\big),$ (10)
is the conditional Shannon entropy.
Lemma 2: Entropy Decomposition. Next, we decompose the total entropy of $\mathbf{TOK}$ to better understand the relationships between conditional entropy and mutual information. According to the entropy decomposition rule [104]:
$H\big(\mathbf{TOK}\big)=H\big(\mathbf{TOK}\mid\mathbf{S}\big)+I\big(\mathbf{TOK};\mathbf{S}\big),$ (11)
where $I\big(\mathbf{TOK};\mathbf{S}\big)$ denotes the mutual information between $\mathbf{TOK}$ and $\mathbf{S}$. Since entropy terms evolve during training, we can rewrite Equation (11) as:
$\Delta H\big(\mathbf{TOK}\big)=\Delta H\big(\mathbf{TOK}|\mathbf{S}\big)+\Delta I\big(\mathbf{TOK};\mathbf{S}\big).$ (12)
Lemma 3: Information Equivalence During Feature Clustering. Assuming the clustering algorithm uses a sufficiently large centroid number and that $\mathbf{S}$ is highly compressible (since our learning objective minimizes its bitrate), the clustered feature $\mathcal{C}$ retains nearly all the information from $\mathbf{S}$,
$H\big(\mathbf{S}\mid\mathcal{C}\big)\approx 0.$ (13)
The chain rule [105] for mutual information yields
$I\big(\mathbf{TOK};\mathbf{S},\mathcal{C}\big)=I\big(\mathbf{TOK};\mathcal{C}\big)+I\big(\mathbf{TOK};\mathbf{S}\mid\mathcal{C}\big),$ (14)
$I\big(\mathbf{TOK};\mathbf{S},\mathcal{C}\big)=I\big(\mathbf{TOK};\mathbf{S}\big)+I\big(\mathbf{TOK};\mathcal{C}\mid\mathbf{S}\big).$ (15)
Equating Equations (14) and  (15) gives,
$I\big(\mathbf{TOK};\mathbf{S}\big)-I\big(\mathbf{TOK};\mathcal{C}\big)$ (16)
$=\Bigl[I\big(\mathbf{TOK};\mathbf{S}\mid\mathcal{C}\big)-I\big(\mathbf{TOK};\mathcal{C}\mid\mathbf{S}\big)\Bigr].$
According to the entropy decomposition rule [104] and the non-negativity of the entropy, we have,
$0\leq I\big(\mathbf{TOK};\mathbf{S}\mid\mathcal{C}\big)\leq H\big(\mathbf{S}\mid\mathcal{C}\big)\approx 0.$ (17)
Since $\mathcal{C}$ is derived from $\mathbf{S}$, we also have,
$I\big(\mathbf{TOK};\mathcal{C}\mid\mathbf{S}\big)=0.$ (18)
Combining Equation (16), (17), and (18), we obtain
$\left|I\big(\mathbf{TOK};\mathbf{S}\big)-I\big(\mathbf{TOK};\mathcal{C}\big)\right|$ (19)
$=\left|I\big(\mathbf{TOK};\mathbf{S}\mid\mathcal{C}\big)-I\big(\mathbf{TOK};\mathcal{C}\mid\mathbf{S}\big)\right|\approx 0.$
This indicates that the mutual information between $\mathbf{TOK}$ and $\mathcal{C}$ is very close to that between $\mathbf{TOK}$ and $\mathbf{S}$,
$I\big(\mathbf{TOK};\mathbf{S}\big)\approx I\big(\mathbf{TOK};\mathcal{C}\big).$ (20)
Theorem 1: NSS Loss Increases the Semantic Separability. Now, building upon the previous lemmas, we demonstrate how the NSS loss enhances the video token semantic separability. Recall that the NSS loss in Equation (5) can be reformulated as,
$\mathcal{L}_{Sem}=\beta\underbrace{\big[-\mathbb{E}_{P_{r}}\,\log P(\mathbf{TOK}\mid\mathbf{S})\big]}_{term~1}+\mathcal{L}_{MAE},$ (21)
where $\beta$ is a balancing parameter.
  • The term 1 is an upper bound on $H\big(\mathbf{TOK}\mid\mathbf{S}\big)$, from Equation (9) and (10). Minimizing it reduces $H\big(\mathbf{TOK}\mid\mathbf{S}\big)$, i.e., $\Delta H\big(\mathbf{TOK}\mid\mathbf{S}\big)<0$.
  • The term $\mathcal{L}_{MAE}$ encourages highly semantic token representations such as local concepts, intuitively reducing the overall entropy, i.e., $\Delta H\big(\mathbf{TOK}\big)<0$.
By choosing appropriate $\beta$, $\Delta H\big(\mathbf{TOK}\mid\mathbf{S}\big)$ can be decreased more quickly than $\Delta H\big(\mathbf{TOK}\big)$, while keeping that $\mathcal{L}_{MAE}$ still converges. From Equation (12), this achieves,
$\Delta I\big(\mathbf{TOK};\mathbf{S}\big)=\Delta H\big(\mathbf{TOK}\big)-\Delta H\big(\mathbf{TOK}\mid\mathbf{S}\big)>0,$ (22)
With Equation (20), we have,
$\Delta I\bigl(\mathbf{TOK};\mathcal{C}\bigr)>0,$ (23)
where $\mathcal{C}$ denotes discrete clustered label set.
Then, considering that the variational bounds [106] of mutual information between the video patch token feature and the discrete label can be given as,
$I\big(\mathbf{TOK};\mathcal{C}\big)$ (24)
$\leq\mathbb{E}\!\left[\frac{1}{N}\sum_{i=1}^{N}\log\frac{P_{r}\big(\mathcal{C}[i]\mid{Tok}[i]\big)}{\frac{1}{N-1}\sum_{k\neq i}P_{r}\big(\mathcal{C}[i]\mid{Tok}[k]\big)}\right],$
where $\mathcal{C}[i]$ and $Tok[i]$ are at the same spatial position in the same video. $N$ is the number of all video patch tokens within the training batch.
Combining Equations (23) and (24), we conclude that the NSS objective increases the mutual information between the token and clustered semantic features, thereby enhancing the semantic separability of the token features.
Refer to caption Refer to caption
(a) (b)
Visualization. To provide a more intuitive understanding how NSS loss improves the semantic boundaries, we further visualize the patch features. First, we randomly select one hundred videos from the UCF101 dataset. Then, we cluster the patch features of these videos into individual groups using the K-Means algorithm [107]. The cluster ID is then utilized as the pseudo-category for each patch. Subsequently, the features are visualized using the t-SNE technique [108]. As compared in Figure 3, although the vanilla MAE loss can learn roughly separable semantic representation, our non-semantics suppressed loss $\mathcal{L}_{Sem}$ learns much clearer semantic boundaries, indicating less semantic noises. This also saves the semantic bitcost (0.0074bpp v.s. 0.0028bpp for $\mathcal{L}_{MAE}$ and $\mathcal{L}_{Sem}$, respectively).